OpenStudy (anonymous):

Which of the following is equal to the square root of the cube root of 5?

OpenStudy (anonymous):



OpenStudy (freckles):

\[\sqrt{ \sqrt[3]{5}}=\sqrt{5^\frac{1}{3}} =(5^\frac{1}{3} )^\frac{1}{2}\]

OpenStudy (freckles):

use \[(x^r)^s=x^{r \cdot s}\]

OpenStudy (freckles):

this is a law of exponent

OpenStudy (anonymous):

@freckles so do we add the fractions together

OpenStudy (freckles):

no r*s means multiply r and s